OpenStudy (anonymous):

Write an equation and solve it to find the length of the longer side of the rectangle. 2x + 10 = 125; x = 57.5 cm x + 5 = 125; x = 120 cm 5x = 125; x = 25 cm 25x = 125; x = 5 cm

OpenStudy (agent0smith):

Is there a pic? Which is the longer side?

OpenStudy (anonymous):

hold on ill show you

OpenStudy (anonymous):

The rectangle below has a total perimeter of 125 cm.

OpenStudy (agent0smith):

 call the other side x. Perimeter is all the sides added together, and we know it equals 125: x+x+5+5 = 125

OpenStudy (anonymous):

okay ..hm

OpenStudy (agent0smith):

Make sense? Now just simplify that equation. Combine like terms.

OpenStudy (anonymous):

i need help doing that.

OpenStudy (agent0smith):

What's x+x? remember that it's really 1x... so 1x+1x = ?

OpenStudy (anonymous):

2x

OpenStudy (agent0smith):

Yep. So we have 2x + 5 + 5 = 125 and now just add the 5s together.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

??

OpenStudy (agent0smith):

i know you can do this part... 5+5 = ? so 2x + 5 + 5 = 125 when you add the 5's becomes...

OpenStudy (anonymous):

10

OpenStudy (agent0smith):

so now it's 2x + 10 = 125 all we did was combine the x's and the 5's in x + x + 5 + 5 = 125

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

@agent0smith can you please help me with my question.

OpenStudy (anonymous):

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below where A = the pounds of Arabica and R = pounds of Robusta. 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds

OpenStudy (agent0smith):

1000 pounds of arabica, which is A... so A=1000 700A + 1200R = 1,000,000 so plug that in here... replace A with 1000: 700*1000 + 1200R = 1,000,000 700*1000 is 700,000 so 700,000 + 1200R = 1,000,000 subtract 700,000 on both sides 1200R = 1,000,000 - 700,000 1200R = 300,000 Now just divide both sides by 1200, that'll give R.

OpenStudy (anonymous):

oh okay cool

OpenStudy (agent0smith):

Did you find R?

OpenStudy (anonymous):

250?

OpenStudy (agent0smith):

Yes, good job :)

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

Evan is making a table that will be created in the shape of the figure below. The table top is a triangle attached to a rectangle. To purchase the right amount of paint, he needs to know the area of the table top. He can only spend $10 on paint, which is enough to cover 150 ft2 of surface area. What is the maximum length of the base of the rectangle he can build?

OpenStudy (anonymous):